OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

hero (hero):

Using y = mx + b only: 10 = m + b 1 = -2m + b b = b 10 - m = 2m + 1 10 - 1 = 2m + 1 9 = 3m 9/3 =m 3 = m 10 - 3 = b 7 = b Therefore: y = 3x + 7

OpenStudy (anonymous):

thats not one of the choices

OpenStudy (anonymous):

First you need to find the slope of the line, or m. The slope is the rise over run. For two points this is equivalent to. \[\frac{ y _{1} - y _{2} }{ x _{1} - _{2} }\] In this case, the slope, or m, is.\[m = \frac{ 1 - 10 }{ -2 - 1} = \frac{ -9 }{ -3 } = 3\] Now we need to find what the y intercept, or b, is. The y intercept is what point on the y-axis the line intersects at. We can calculate this via the following equation.\[y = mx + b\] What we must do is to substitute the slope, m, and one of the points, y and x, with their respective values.\[10 = 3(1) + b\]\[7 = b\] Now that we know the slope and the y intercept we can substitute that into the formula to get our answer of...\[y = 3x + 7\]